# Reducing the Scope of Language Models

### v2

@@ 4.3 Accepting multiple tasks
4.4 Representation Analysis
To get a sense for how different methods, operate, we study how the representations change before and after training on Mistral-7b-Instruct-v0.2. In particular we look at cosine distance between the original representation of a token, and the same token after training. If there are patterns in these changes, it should give us some indication as to how different methods operate. In Figure 4 we show that DPO and SFT only change the representations of the tail of the context. Hence it makes sense why CB is more robust under attack (explored in Zou et al. (2024) and Appendix A.3): all representations have changed, so it is difficult to find a way to circumvent the changed behavior, while DPO and SFT have ”cracks” which can be exploited.
@@ 4.4 Representation Analysis
The effect is particularly clear on the in-distribution rejection set, but preceding sections demonstrated that most methods were fairly comparable in distribution. Out of distribution, the effect of CB is much less, though still there is a much more substantial difference from the original model than SFT or DPO which make only small changes to the tail of context in deeper layers. With SFT-CB, we can clearly see the layering of the tail edit as well as the orthogonalization across the entire context.
4.5 Additional analysis
@@
Adversarial Evaluation: We examine the behavior of different methods under adversarial prompts in Appendix A.3, where we find that CB-based methods are more robust than others, echoing results of Zou et al. (2024).
Model Scale: In Appendix A.2 we examine scoping behavior with different model scales, finding consistent results across scales, except for Probe which struggles with smaller models.
@@ 5 Discussion
Though current language models are generally applicable, there is still a need at deployment time to specify the kinds of queries they should and should not be able to answer. Otherwise, agents deployed in the wild may be easy to distract, and if used as a part of a pipeline may lead to cascading errors. Scoping is crucial. In this work, we conducted a comprehensive empirical study of scoping language models using three model families and multiple tasks.
The general takeaways are many, but firstly we find across many cases that system prompting is very insufficient. Different methods had varying performance on different model and dataset combinations, but by and large we found that SFT works well when the training data is diverse, and CB works well when it is not and there is very little data, likely as the orthogonalization objective is easier to optimize. Probe is also a good choice as long as the probe is strong enough to disentangle the model representations, but requires additional inference time due to the extra model. Layering SFT and CB also generally combines the performance of both, though if either SFT or CB fails then the layering will suffer too.
We saw that it was possible to scope across language model scales, for multiple tasks at a time and for very fine-grained tasks. We demonstrated that different methods have different effects on the internal representations of the models: SFT and DPO only modify the tail of the language model context, unlike CB which modifies representations across the context. Such different behavior may explain why CB is stronger under adversarial attacks (Appendix A.3), the original setting it was proposed for (Zou et al., 2024).
Though all of this suggests a favorable story about CB, it is quite difficult to tune. Performance can fluctuate wildly based on very small step-count differences, and it can be important to tune target layers on a per-model basis. Such considerations lead us to make the generic recommendation that, except when data diversity is low, it may be wise to stick to simple methods like SFT and probing.
@@ A.1 Data diversity across different tasks
We continue discussion on data diversity from Section 4.2. Here we examine performance with varying rejection set diversity using Mistral-7b-Instruct-v0.2 on more tasks than just Sentiment Analysis.
@@
A.2 Scoping over different scales
@@
Here we ask: how does scoping vary with model scale? We use the Llama-3.1 family (Grattafiori et al., 2024) as an example. We fix the accept task as Sentiment Analysis again and check how different methods behave with 1B, 3B and 8B parameter instruct-models. We see in Figure 6 broadly that larger models lead to improved results. The only other major surprise is that probing performs poorly with smaller language models. This may be due to the fact that it is difficult for a small probe to disentangle the representations of smaller models, as they are lower dimension and the information is more compressed.
A.3 Robustness to Adversarial Prompts
The original CB methodology (Zou et al., 2024), and many related works discussed, focus on robustness to adversarial prompts. If models are to be deployed, we might expect that users could attempt adversarial attacks against the deployment. Here we ask: if models are scoped, how robust are they to adversarial prompts? We take Mistral-7b-Instruct-v0.2 as a case study here.
Our threat model is of a black-box access. In particular, we assume that the users are allowed to edit only the instruction text, and do not even have access to modifying the system prompt, which would be true for text-based API access. We implement and test a number of different black-box adversarial attacks:
@@ A.3 Robustness to Adversarial Prompts
Base-64 translation (b64): Following Wei et al. (2023), we translate instructions into base-64, then prompt the language model. After receiving the response if it is valid base-64 (which is very often), we translate it back to text.
Few-shot prompting (Few-shot): We provide a few-shot example from the evaluation set, where we draw a training query and completion and then prompt the next round with a new query. This is similar to the Many-Shot attack explored by  Anil et al. (2024).
@@
Prefill: In this attack, append a generic prefilling output (”Sure, here’s a response to your request:”), to the end of the user instruction. This follows common practice and has been a strong attack (Wei et al., 2023; Zou et al., 2024).
Tree of Attacks with Pruning (TAP): TAP (Mehrotra et al., 2023) is an adversarial prompting method whereby an attacking language model iteratively attempts to jailbreak a target language model (here our model to evaluate). TAP uses an actor-critic loop to rewrite prompts based on whether the language model was jailbroken in the previous step, and explores a tree to find the best performing prompt. For more details and hyperparameters, see Appendix B. It is a strong black-box optimization-based jailbreaking method, and was among the strongest attacks for CB in the original setting (Zou et al., 2024). As TAP is quite expensive to run, we only test 10 prompts per dataset.
@@
Takeaways: Notably, both Sys. and SFT are quite poor. DPO as well has many issues. Probe is quite strong, but also has a tendency to reject accept tasks, which is undesirable. Thus it appears CB, and SFT-CB, strike a nice balance between in and out of distribution rejection, as well as letting desired prompts pass through. In all of these evaluations it is clear that none of these methods are even close to perfect, so there is still much work to be done. Such results are quite distinct from the safety picture presented by Zou et al. (2024), perhaps as the domains are not quite as simple as safe vs. unsafe prompts. Still, the spirit of the results in Zou et al. (2024) appear to be true: CB seems more robust to adversarial attacks than baselines, with the exception of the Probe which tends to reject even on accept tasks. One additional point on the b64 attack, which appears to bypass all models: the completions tend to either be generic base-64 encoded response (e.g. ”Hello world!”), or invalid base-64.
A.4 Precise Scoping
@@
Sentiment Analysis: For task performance, unsurprisingly SFT-based methods are best. Strangely DPO seems to suffer when SA-FR is included in the rejection set. All methods have no rejections on the accept task. For the fine-grained rejection set, all methods do well (except Sys.) when it is included in the rejection set, but CB-based methods do best when it is not (see last column). On in-distribution rejection, all methods do well. For out of distribution, we see that CB, SFT-CB and Probe are best on the low-diversity case (only SA-FR), while as the distribution expands other methods catch up echoing results in Section 4.2.
@@
A.5 Ablations on Circuit Breaker hyperparameters
@@
A.6 Effect of Data Quantity
@@
Sentiment Analysis: Perhaps unsurprisingly, SFT-based methods are best across the board. Interestingly, very little data is needed for this task and scores are roughly flat. On the accept set, rejection rates are also flat with the number of prompts, and the Probe always rejects a large number. In-distribution, the major trned to note is that both DPO and Probe are quite stable and strong across number of prompts, but CB appears quite unstable and seesaws. This may be due to difficulty optimizing for orthogonality. A similar trend is visible in the OOD case.
@@
A.7 Effect of LoRA Rank
@@ B.1 Training
For CB, we follow Zou et al. (2024) and add LoRA’s to all matrix parameters (both Attention and MLP), but only on the first 20 layers. We use layers 10 and 20 as targets for the representations, as described by Zou et al. (2024). For SFT and DPO we add LoRA’s to all matrix parameters for all 32 layers. For all experiments we use Adam (Kingma, 2014) without weight decay.
@@ B.2 Evaluation
For every category, when we evaluate we use 256 heldout prompts. The only exception is for TAP (Mehrotra et al., 2023), where due to high costs we only evaluate 10 prompts per category.
@@
Ideally one might choose to use a language model judge for detecting rejection (Zheng et al., 2023). However, given the large number of experiments and evaluations in this work, we found it prohibitively expensive to run all the evaluations through a state-of-the-art API judge. We experimented with using locally hosted language models as judges, with the largest being Llama-3-70B-Instruct (Dubey et al., 2024), but found such detection to have much poorer performance both in precision and recall than the methods described below.
@@
As different methods behave differently, we employ different ways to detect rejection. For all methods besides probing, as the system prompt and tuning will instruct models to respond “I cannot answer that.”, we catch rejection by string matching for a few different tokens that are synonyms for “cannot” at the beginning of the generation. The reason we only match strings early in the generation is that it is possible to switch from reject to accept behavior, so we would like to catch that rejection early on. We do not match strings later on as in practice we never observed language models switching from accept to reject behavior midway through generation and keywords can be used as a part of a legitimate response later in generation. On a sample of 30 completions from accept, reject, and OOD reject sets, we tuned the threshold that such a detector had perfect agreement with manual judgment. This was inspired by common string-based detectors like the one proposed by Zou et al. (2023b) and used by Zeng et al. (2024); Zou et al. (2024).
@@
The full list of strings that we match for the string-based detector is: [”cannot”, ”unable”, ”sorry”, ”can’t”, ”will not”, ”won’t”, ”do not”, ”don’t”, ”not able”]. We arrived at this list after manual verification of a subset of 90 prompts, and we find the performance to match human judgments. As we are tuning, we are unable to simply reuse the list from Zou et al. (2023b) as model completions may change. If we detect any of these words within the first 30 characters (the beginning of the completion), we consider the model to reject that prompt.
@@
For the CB rejection detection, generations break often by repeating the same token over and over again. As exact-matching does not detect such patterns, in addition to string matching described above which will activate when the system prompt is followed, we catch rejection by the existence of a repeated pattern of 4 or more strings within the response. Again on a sample of 30 completions form accept, reject and OOD reject sets, we tuned the threshold for this detector such that it achieved 1 false negative and 0 false positives out of 90 completions. The single false negative was due to to a broken generation of punctuation characters that lacked repetitions. See Appendix C for sample outputs.
@@ B.3 Scoping language models
In Section 4.1, we use Mistral-7B-Instruct-v0.2, Granite-7b-Instruct and Llama-3.1-8B-Instruct for results. Our accept set is SA while the reject set is S, TLD, TC, SC.

### v3

@@ Methods
4 Experiments
All experiments contain evaluations of task performance (Accept Score) on the accept set (which should be high), rejection rate on the in-distribution accept (Accept) set (which should be low) as well as rejection rate on the in-distribution reject set (ID Reject) and out of distribution data (OOD Reject) (which should be high). We describe experiments in broad strokes, and defer precise details on hyperparameters to Appendix B. We begin by presenting results across a variety of language models (Mistral-7B-Instruct-v0.2, granite-7b-instruct, Llama-3.1-8B-Instruct(jiang2023mistral; sudalairaj2024lab; grattafiori2024llama). We explore scale on the Llama-3.2-instruct family, and for other results choose Mistral-7b-instruct as we do not have the compute or space to run every experiment on every model.
Scoping language models
We start with the basic question: is it possible to scope language models? We explore each method across a variety of models using Sentiment Analysis (SA) as our accept task in Figure 13, where we see that although system prompting is insufficient, a broad variety of different methods are successful to different extents with different models.
Rejection set diversity
One of the most critical questions when attempting to restrict the generations of language models is what data might be necessary to do so. If models overfit to a particular data distribution, then it may be difficult to reject requests that were not specified in the training distribution. Thus, here we ask: how much data diversity is necessary in the rejection set to robustly scope models? If very little diversity is needed, and rejection extends to OOD requests, then adapting models to new deployments becomes quite inexpensive. In Figure 14 we study Sentiment Analysis across different models with more and more diverse rejections sets consisting of an increasing number of tasks, and defer results on additional tasks with Mistral-7b-instruct to Appendix A.1.
Accepting multiple tasks
Here we ask: is it possible to still reject tasks when there are multiple tasks in the accept set? Such a setting is natural as most language models will have a few different specific uses, like a programming bot that can write code and also answer questions about documentation. We demonstrate results on Mistral-7b-Instruct-v0.2 in Table 2 with two choices of accepts sets: Classification and Generation (SA,S,TLD,SC,TC,DG) and Math and Programming (PE,GSM8k). We use the opposite set as the ID reject set in this case.
@@ 4 Experiments
Representation Analysis
To get a sense for how different methods operate, we study how the representations change before and after training on Mistral-7b-Instruct-v0.2. In particular we look at cosine distance between the original representation of a token, and the same token after training. If there are patterns in these changes, it should give us some indication as to how different methods operate. In Figure 4 we show that DPO and SFT only change the representations of the tail of the context. Hence it makes sense why CB is more robust under attack (explored in zou2024improving and Appendix A.4): all representations have changed, so it is difficult to find a way to circumvent the changed behavior, while DPO and SFT have “cracks” which can be exploited.
The effect is particularly clear on the in-distribution rejection set, but preceding sections demonstrate that most methods are fairly comparable in distribution. Out of distribution, the effect of CB is much less, though still there is a much more substantial difference from the original model than SFT or DPO which make only small changes to the tail of context in deeper layers. With SFT-CB, we can clearly see the layering of the tail edit as well as the orthogonalization across the entire context.
Additional analysis
@@
Adversarial Evaluation: We examine the behavior of different methods under adversarial prompts in Appendix A.4, where we find that CB-based methods are more robust than others, echoing results of zou2024improving.
Model Scale: In Appendix A.3, we examine scoping behavior with different model scales, finding consistent results except for Probe which struggles with smaller models.
@@ 5 Discussion
Though current language models are generally applicable, there is still a need at deployment time to specify the kinds of queries they should and should not be able to answer. Otherwise, agents deployed in the wild may be easy to distract, and if used as a part of a pipeline may lead to cascading errors. Hence scoping is crucial. In this work, we conducted a comprehensive empirical study of scoping language models using three model families and multiple tasks.
Our findings reveal several key insights. First, system prompting alone is generally inadequate across a range of models and datasets. While performance varies depending on the specific method, model, and dataset, certain trends emerge. Supervised fine-tuning (SFT) tends to perform well when the training data is diverse, whereas Circuit Breakers (CB) is more effective in low-data regimes with less diversity, likely because the orthogonalization objective it employs is easier to optimize in such settings. Probing methods can also yield strong results, provided that the probe is sufficiently expressive to disentangle the model’s internal representations; however, this approach incurs additional inference overhead due to the use of an auxiliary model. Combining SFT and CB typically results in performance that reflects the strengths of both approaches, but this layered method is sensitive to the failure of either component, which can degrade overall performance.
We saw that it was possible to scope across language model scales, for multiple tasks at a time and for very fine-grained tasks. We demonstrated that different methods have different effects on the internal representations of the models: SFT and DPO only modify the tail of the language model context, unlike CB which modifies representations across the context. Such different behavior may explain why CB is stronger under adversarial attacks (Appendix A.4), the original setting it was proposed for (zou2024improving).
While these results indicate that CB can be a promising approach, it presents considerable challenges in practice. Specifically, its performance can fluctuate wildly based on very small step-count difference. Additionally, optimal target layers may need to be selected on a per-model basis, further complicating its application. Given these sensitivities, we recommend defaulting to simpler methods such as supervised fine-tuning (SFT) or probing, particularly in settings where data diversity is not a limiting factor.
@@ A.1 Data diversity across different tasks
We continue discussion on data diversity. Here we examine performance with varying rejection set diversity using Mistral-7b-Instruct-v0.2 on more tasks than just Sentiment Analysis.
@@
A.2 Additional Classification Baselines
We previously compared the performance of scoping methods against a classifier (Probe) deployed in a two-stage (classify-then-generate) setup. For completeness, we compare the performance of the Probe Classifier against other baselines such as LLM judges based on LLama 3.1 8B instruct and Llama 3.2 70B Instruct deployed as classifiers, as well as a fine-tuned classifier based on a state-of-the-art general purpose text embedding based model - GTE 1.5.  [^4] https://huggingface.co/Alibaba-NLP/gte-base-en-v1.5 We train this model by adding a linear head for binary classification and tune all model parameters.
Method Accept ID Reject OOD Reject
Probe Baseline 0.035 $0.924\pm 0.152$ $0.952\pm 0.107$
Finetuned GTE 0.0 1.0 $0.821\pm 0.355$
Llama 3.1 8B Instruct (Sys) 0.203 $0.770\pm 0.433$ $0.801\pm 0.146$
Llama 3.1 8B Instruct (Judge) 0.082 $0.549\pm 0.162$ $0.700\pm 0.283$
Llama 3.2 70B Instruct (Judge) 0.004 $0.945\pm 0.084$ $0.819\pm 0.358$
Table 3: Classifier baselines
We see that prompted solutions at the same scale as the final model (8B, Sys. and Judge) are substantially worse both ID and OOD than the Probe baseline. We also see that at the 70B scale, the judge is still not better OOD. A large judge would not make sense at deployment time due to the high cost of inference. Thus we believe it is important to explore methods of LLM scoping.
For Finetuned GTE, the results are strong compared to many methods, but worse than the Probe baseline at the largest scale. Considering we will need to run the 8B model anyway, a GTE classifier solution is strictly more computationally inefficient and additionally more complex to manage.
A.3 Scoping over different scales
@@
Here we ask: how does scoping vary with model scale? We use the Llama-3.1 family (grattafiori2024llama) as an example. We fix the accept task as Sentiment Analysis again and check how different methods behave with 1B, 3B and 8B parameter instruct-models. We see in Figure 6 broadly that larger models lead to improved results. The only other major surprise is that probing performs poorly with smaller language models. This may be due to the fact that it is difficult for a small probe to disentangle the representations of smaller models, as they are lower dimension and the information is more compressed.
A.4 Robustness to Adversarial Prompts
The original CB methodology (zou2024improving), and many related works discussed, focus on robustness to adversarial prompts. If models are to be deployed, we might expect that users could attempt adversarial attacks against the deployment. Here we ask: if models are scoped, how robust are they to adversarial prompts? We take Mistral-7b-Instruct-v0.2 as a case study here.
Our threat model is of black-box access. In particular, we assume that the users are allowed to edit only the instruction text, and do not even have access to modifying the system prompt, which would be true for text-based API access. We implement and test a number of different black-box adversarial attacks:
@@ A.4 Robustness to Adversarial Prompts
Base-64 translation (b64): Following wei2024jailbroken, we translate instructions into base-64, then prompt the language model. After receiving the response if it is valid base-64 (which is very often), we translate it back to text.
Few-shot prompting (Few-shot): We provide a few-shot example from the evaluation set, where we draw a training query and completion and then prompt the next round with a new query. This is similar to the Many-Shot attack explored by  anil2024many.
@@
Prefill: In this attack, append a generic prefilling output (”Sure, here’s a response to your request:”), to the end of the user instruction. This follows common practice and has been a strong attack (wei2024jailbroken; zou2024improving).
Tree of Attacks with Pruning (TAP): TAP (mehrotra2023tree) is an adversarial prompting method whereby an attacking language model iteratively attempts to jailbreak a target language model (here our model to evaluate). TAP uses an actor-critic loop to rewrite prompts based on whether the language model was jailbroken in the previous step, and explores a tree to find the best performing prompt. For more details and hyperparameters, see Appendix of Experimental Details. It is a strong black-box optimization-based jailbreaking method, and was among the strongest attacks for CB in the original setting (zou2024improving). As TAP is quite expensive to run, we only test 10 prompts per dataset.
@@
Takeaways: Notably, both Sys. and SFT are quite poor. DPO as well has many issues. Probe is quite strong, but also has a tendency to reject accept tasks, which is undesirable. Thus it appears CB, and SFT-CB, strike a nice balance between in and out of distribution rejection, as well as letting desired prompts pass through. In all of these evaluations it is clear that none of these methods are even close to perfect, so there is still much work to be done. Such results are quite distinct from the safety picture presented by zou2024improving, perhaps as the domains are not quite as simple as safe vs. unsafe prompts. Still, the spirit of the results in zou2024improving appear to be true: CB seems more robust to adversarial attacks than baselines, with the exception of the Probe which tends to reject even on accept tasks. One additional point on the b64 attack, which appears to bypass all models: the completions tend to either be generic base-64 encoded response (e.g. ”Hello world!”), or invalid base-64.
A.5 Precise Scoping
@@
Sentiment Analysis: For task performance, unsurprisingly SFT-based methods are best. Strangely DPO seems to suffer when SA-FR is included in the rejection set. All methods have no rejections on the accept task. For the fine-grained rejection set, all methods do well (except Sys.) when it is included in the rejection set, but CB-based methods do best when it is not (see last column). On in-distribution rejection, all methods do well. For out of distribution, we see that CB, SFT-CB and Probe are best on the low-diversity case (only SA-FR), while as the distribution expands other methods catch up, echoing previous results.
@@
A.6 Ablations on Circuit Breaker hyperparameters
@@
A.7 Effect of Data Quantity
@@
Sentiment Analysis: Perhaps unsurprisingly, SFT-based methods are best across the board. Interestingly, very little data is needed for this task and scores are roughly flat. On the accept set, rejection rates are also flat with the number of prompts, and the Probe always rejects a large number. In-distribution, the major trend to note is that both DPO and Probe are quite stable and strong across number of prompts, but CB appears quite unstable and seesaws. This may be due to difficulty optimizing for orthogonality. A similar trend is visible in the OOD case.
@@
A.8 Effect of LoRA Rank
@@ B.1 Training
For CB, we follow zou2024improving and add LoRA to all matrix parameters (both Attention and MLP), but only on the first 20 layers. We use layers 10 and 20 as targets for the representations, as described by zou2024improving. For SFT and DPO we add LoRA to all matrix parameters for all 32 layers. For all experiments we use Adam (kingma2014adam) without weight decay.
@@ B.2 Evaluation
For every category, when we evaluate we use 256 heldout prompts. The only exception is for TAP (mehrotra2023tree), where due to high costs we only evaluate 10 prompts per category.
@@
Ideally one might choose to use a language model judge for detecting rejection (zheng2023judging). However, given the large number of experiments and evaluations in this work, we found it prohibitively expensive to run all the evaluations through a state-of-the-art API judge. We experimented with using locally hosted language models as judges, with the largest being Llama-3-70B-Instruct (dubey2024llama), but found such detection to have much poorer performance both in precision and recall than the methods described below.
@@
As different methods behave differently, we employ different ways to detect rejection. For all methods besides probing, as the system prompt and tuning will instruct models to respond “I cannot answer that.”, we catch rejection by string matching for a few different tokens that are synonyms for “cannot” at the beginning of the generation. The reason we only match strings early in the generation is that it is possible to switch from reject to accept behavior, so we would like to catch that rejection early on. We do not match strings later on as in practice we never observed language models switching from accept to reject behavior midway through generation and keywords can be used as a part of a legitimate response later in generation. On a sample of 30 completions from accept, reject, and OOD reject sets, we tuned the threshold that such a detector had perfect agreement with manual judgment. This was inspired by common string-based detectors like the one proposed by zou2023universal and used by zeng2024johnny; zou2024improving.
@@
The full list of strings that we match for the string-based detector is: [”cannot”, ”unable”, ”sorry”, ”can’t”, ”will not”, ”won’t”, ”do not”, ”don’t”, ”not able”]. We arrived at this list after manual verification of a subset of 90 prompts, and we find the performance to match human judgments. As we are tuning, we are unable to simply reuse the list from zou2023universal as model completions may change. If we detect any of these words within the first 30 characters (the beginning of the completion), we consider the model to reject that prompt.
@@
For the CB rejection detection, generations break often by repeating the same token over and over again. As exact-matching does not detect such patterns, in addition to string matching described above which will activate when the system prompt is followed, we catch rejection by the existence of a repeated pattern of 4 or more strings within the response. Again on a sample of 30 completions form accept, reject and OOD reject sets, we tuned the threshold for this detector such that it achieved 1 false negative and 0 false positives out of 90 completions. The single false negative was due to to a broken generation of punctuation characters that lacked repetitions. See Appendix of Sample outputs.
@@ B.3 Scoping language models
In Section of Scoping language models, we use Mistral-7B-Instruct-v0.2, Granite-7b-Instruct and Llama-3.1-8B-Instruct for results. Our accept set is SA while the reject set is S, TLD, TC, SC.
@@ B.4 Rejection Set Diversity
(a) Mistral-7B-Instruct
(b) Llama-3.1-8B-instruct
(c) Granite-7B-instruct